OpenStudy (anonymous):

Alg2. find the coordinates of where the graph intersects, show substitution of both coordinates x-2y=-4 3x+2y=12

OpenStudy (anonymous):

Add first and second equations 4x=8 x=8/4=2, plug x=2 into 1st eq. 2-2y=-4 2+4=2y 2y=6 y=6/2=3 intersection (2,3) To check plug into both eq-ns

OpenStudy (anonymous):

1st 2-2*3=-4 True 2nd 3*2+2*3=12 True